OpenStudy (anonymous):

i am unable to do higher calculations ..the problems seems easy but reaching the final answer seems difficult ...i get confused what to do

OpenStudy (anonymous):

what is the problem? what kind of higher calculations?

OpenStudy (anonymous):

at the time when it cums to shifting decimal i totally get confused and wonder what to multiply with which number

OpenStudy (anonymous):

\[6.67x10^{-11}\times7.36\times10^{22}\times6\times10^{24}\div(3.8\times3.8\times10^{5}\times10^{5}\]

OpenStudy (anonymous):

how to solve these type of questions fast which i am unable to do

OpenStudy (turingtest):

just multiply all the 10^x stuff first, i.e.\[{6.67\times10^{-11}\times7.36\times10^{22}\times6\times10^{24}\over3.8\times3.8\times10^{5}\times10^{5}}\\={6.67\times7.36\times6\over3.8\times3.8}\times{10^{-11}\times10^{22}\times10^{24}\over10^5\times10^5}\\={6.67\times7.36\times6\over3.8\times3.8}\times10^{25}\]

OpenStudy (anonymous):

just for the 10^ part, for their exponents, which is in numerator, it's +, which is in denominator, it's - . Therefore, the exponent number is 11+22+24 -5-5= 47. Now , position it as it is, you have \(\Large 10^{47}\)

OpenStudy (turingtest):

the first exponent was -11, so it winds up as \(\cdots\times10^{25}\)

OpenStudy (anonymous):

ohh, I am sorry, my computer didn't show the minus sign of it.

OpenStudy (kainui):

If you ever forget how exponents add, just do something with x^3 and x^2. \[x^3*x^2=(x*x*x)*(x*x)=x^5=x^{3+2}\] Do the same thing with division, figure it out on your own, it's almost identical to this. Don't lose sight of the idea that all an exponent is is simply notation to tell you how many times to multiply a number by itself. (Most of the time... lol)

OpenStudy (chenping):

just press calculator dude :D

OpenStudy (anonymous):

unfortunately its not allowed in d exams :(

OpenStudy (rajat97):

you can use a log book

OpenStudy (chenping):

hahaha, Turing has given the ans alrdy.. also, rmb that we dun giv ans, we shud giv solution instead :)

OpenStudy (anonymous):

Do 10 powers seperately and then the significant numbers for example calculate \[\frac{2 \times 10^{-13} \times -5 \times 10^{+21} }{6 \times 10^{13} \times 3 \times 10^{5}}\]

OpenStudy (anonymous):

can u try this?

OpenStudy (anonymous):

\[-10^{-27}\div12\]

OpenStudy (anonymous):

correct ??

OpenStudy (anonymous):

wait :P

OpenStudy (anonymous):

first bring everything to numerator \[10^{-13} \times 10^{-13} \times 10^{21} \times 10^{-5}\]

OpenStudy (anonymous):

is that ok? makes sense? when u bring denominator exponents to numerator their signs flip! yeaa?

OpenStudy (anonymous):

yaa i got that but was my answer correct ??

OpenStudy (anonymous):

no it wasn't.. now try again.. tell me what is the answer?

OpenStudy (anonymous):

\[0.83\times10^{-8}\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

nope.. next step is you write it as \[10 ^{-13-13+21-5}\] lets only talk about 10 powers.. forget about the numbers ok.. SO now tell.. whats the ten power gonna be?

OpenStudy (anonymous):

its 10 TO THe PowER -8 again d same thing

OpenStudy (anonymous):

-13-13+21-5 = -26+21-5 try again ? :P..

OpenStudy (anonymous):

omg i thought thats 12 instead of 13

OpenStudy (anonymous):

:P :P

OpenStudy (anonymous):

ok so now whats the answer?

OpenStudy (anonymous):

its 10 to the power -10

OpenStudy (anonymous):

thats it.. so now comes a bonus question for you \[\frac{3 \times 10^{-23} \times -4 \times 10^{21} }{8 \times 10^{-42} \times 6 \times 10^{14}}\] ............ 2 points..2 tries.. give me complete answer.. ready? go!

OpenStudy (anonymous):

-1/4 times 10 to the power 26

OpenStudy (chenping):

Bullseye! Congratz! :)

OpenStudy (anonymous):

correct ?

OpenStudy (anonymous):

yuhu!! :P

OpenStudy (anonymous):

yea.. u got that right.. u passed my crash course :P..